# On the Value of Tokeniser Pretraining in Physics Foundation Models

### 05 Mar 2026 v1

@@ 4.1 Key Findings
This work provides the first systematic study of tokeniser pretraining for physics foundation models. We demonstrate that pretraining substantially accelerates downstream training under limited computational budgets, with benefits critically dependent on domain alignment between pretraining and downstream tasks. Key findings include:
@@
Scale-dependent learning dynamics. Pretraining affects learning differently across spatial scales (Figure 2). At low frequencies, in-domain models achieve excellent final performance (NEPS $\sim$0.001) compared to models without pretraining (NEPS $\sim$0.02). At mid frequencies, in-domain models maintain consistently lower error throughout training (NEPS $\sim$0.1–0.3 vs $\sim$0.3–1.0 without pretraining). At high frequencies, all models show poor quality (NEPS $\approx$ 1.0), but in-domain pretraining enables continuous improvement whereas models without pretraining show progressive degradation.
@@ 4.2 Limitations
Several limitations warrant discussion. First, out-of-domain pretraining provides modest benefits, and freezing the tokeniser in this setting degrades performance below the no-pretraining baseline. One explanation is limited field overlap between pretraining and downstream datasets: only one of four physical quantities in the Euler dataset (Table 1) appears in the multi-dataset mixture, while the others are encountered only during downstream training. Future work could investigate whether greater field overlap improves out-of-domain transfer, or develop methods to help pretrained tokenisers adapt to previously unseen physical fields.
@@ 4.3 Future Directions
Downstream task diversity. While this work focuses on autoregressive rollout prediction, pretrained tokenisers may prove valuable across broader physics emulation tasks. Evaluating the same pretrained tokeniser on diverse downstream objectives, such as masked token modelling, physical parameter inference, or inverse problems, would reveal which tasks benefit most from autoencoding pretraining.
Tokeniser architecture & training improvements. Investigating tokeniser design could yield more general insights and improve both pretraining and downstream accuracy. Promising directions include: evaluating alternative architectures to identify inductive biases better suited to physics data; scaling studies to determine whether benefits compound at larger capacity and with more diverse pretraining data; comparing training objectives (reconstruction, perceptual, adversarial losses) and quantisation strategies; and determining optimal pretraining budgets (measured in training steps, data fraction, or target accuracy) to maximise downstream performance while minimising computational cost.
Data composition & domain alignment. Our results show that domain alignment between pretraining and downstream tasks significantly impacts performance gains, raising an important question about data composition: given a fixed pretraining budget, does training on diverse multi-physics mixtures produce more versatile and generalisable tokenisers than domain-specific pretraining? Furthermore, systematically varying domain similarity, from small differences (boundary conditions, physical parameters) to large ones (governing equations), would quantify how domain proximity impacts transfer. Additionally, investigating how pretraining benefits depend on downstream data size would clarify when tokeniser pretraining justifies its computational cost and how data scarcity affects these gains.
@@
Together, these research directions would yield principled guidelines for tokeniser pretraining in physics foundation models. If future investigations demonstrate consistent out-of-domain benefits, they would justify developing and releasing high-quality pretrained tokenisers as community resources across scientific and engineering domains.
@@ 5 Conclusion
This work provides the first systematic investigation of tokeniser pretraining for physics foundation models. We demonstrate that pretraining offers a practical path to improved training efficiency, with benefits that depend critically on domain alignment between pretraining and downstream tasks. We also find that freezing most of an in-domain pretrained tokeniser can improve long-horizon rollouts, suggesting a simple regularisation strategy for stable autoregressive prediction. These findings position tokeniser pretraining as a simple yet effective mechanism for improving the training of physics foundation models.
@@ Appendix A Dataset Properties
Evolution of different gases starting with piecewise constant initial data in quadrants. The evolution can give rise to shocks, rarefaction waves, contact discontinuities, interaction with each other and domain walls. The evolved fields are density $\rho$, energy $e$, pressure $p$, and velocity components $u$ and $v$, and the system follows Euler equations for a compressible gas. The fields are sampled on a uniform $\times512$ Cartesian grid over 100 time steps. The dataset comprises 20 parameter configurations, varying in boundary condition type (open or periodic) and gas constant $\gamma$, with 500 trajectories per configuration.
@@
Simulation of a continuum model for biological active matter, comprising rod-like particles immersed in a Stokes fluid. The evolved fields include a tracer-like concentration field $c$, fluid velocity $(u,v)$, an orientation tensor (describing alignment of the active agents), and a strain-rate tensor. The data is sampled on a uniform $\times256$ Cartesian grid over 81 time steps, across 45 parameter sets and 5 trajectories per parameter set.
@@ Appendix B Tokeniser Architecture and Flexible Compression
MAGVIT-Simple autoencodes temporal sequences of 2D or 3D simulations ${\mathbf{x}}\in{\mathbb{R}}^{T\times S_{1}\times\dots\times S_{k}}$, where $k\in\{2,3\}$. The temporal size $T$ is set to ”1 + a power of two”, and the spatial sizes $S_{i}$ are powers of two. Figure 4 illustrates the tokeniser architecture.
We employ causal convolutions, as introduced in MAGVIT-2, which enable joint image-video tokenisation and have been shown to outperform non-causal convolutions and attention-based blocks Yu et al. (2024): for 2D simulations, a 3D causal convolutional layer with kernel size $(k_{t},k_{s},k_{s})$ is applied to the input [^3] Operating on 3D simulations would require using 4D causal convolutions.. Spatial padding is applied symmetrically to both ends of each spatial dimension and is configured to preserve the input’s spatial size. Temporal padding is applied only at the beginning (i.e., $k_{t}-1$ leading frames), with no padding at the end. This preserves the temporal size of the input and ensures that each output timeframe depends only on the current and preceding frames. Padding values are set to zero.
The encoder and decoder heads use adaptive causal 3D convolutions that support a variable number of input physical fields. A single convolutional kernel is learned for the maximum number of fields; at runtime, rows and columns corresponding to active fields are selected from the weight matrix, and the output is rescaled by $\sqrt{C_{\text{total}}/C_{\text{active}}}$ to compensate for reduced fan-in. This follows the field embedding approach from MPP (McCabe et al., 2024).
@@
Given a scaling pair $(s_{t},s_{s})$, we define the spatial and temporal kernel sizes as
$k_{s}=s_{s}\qquad\qquad\qquad k_{t}=\begin{cases}1+s_{t},&\text{if }s_{t}>1,\\ s_{t},&\text{if }s_{t}=1.\end{cases}$ (6)
@@
pixel in the output - a strategy we adopt in our flexible depth-to-space layers. Finally, because this operation introduces additional time frames, we follow MAGVIT-2 (Yu et al., 2024) and discard the first $s_{t}-1$ frames from the output.
@@ Appendix C Detailed Training Recipe
This appendix provides the complete set of hyperparameters for both training stages. High-level training details are described in Subsection 2.5, and the tokeniser architecture is described in Appendix B.
Table 6 summarises the training hyperparameters for both stages.
@@
For rollout training, we use an inverse-square-root schedule with linear warmup and square-root cooldown, stepping once per epoch. Let $e$ denote the current epoch (1-indexed), with total epochs $E=100$, warmup duration $W=10$, cooldown duration $C=10$, warmup starting factor $\alpha=0.1$, and decay offset $\kappa=128$. The schedule has three phases:
@@
$\text{lr}(e)=\text{lr}_{0}\left(\alpha+(1-\alpha)\frac{e}{W}\right).$ (7)
@@
$\text{lr}(e)=\frac{\text{lr}_{0}}{1+\sqrt{s+\kappa}-\sqrt{\kappa}}\,$, (8)
where s=eWs=e-W is the number of epochs since warmup ended.
Cooldown (e>ECe>E-C): The learning rate ramps down to zero with a square-root profile:
$\text{lr}(e)=\text{lr}_{\text{end}}\left(1-\sqrt{\frac{t}{C}}\right),$ (9)
where t=e(EC)t=e-(E-C) counts epochs into cooldown and $\text{lr}_{\text{end}}$ is the learning rate at the end of decay.
@@
Each GPU processes a disjoint shard of the training data. For tokeniser pretraining, each shard contains 21,000 unique batches (of size 2). Once all unique batches have been consumed, the data loader reseeds its random number generator and begins iterating over the dataset again, allowing resampling in subsequent passes. For rollout training, the same mechanism applies but each shard contains 2,100 unique batches.
@@
In downsampling layers, stride equals the chosen compression $(s_{t},s_{s},s_{s})$ (see Table 7). The convolution kernel size is tied to the stride: temporal size is $k_{t}=1+s_{t}$ when $s_{t}>1$ and $k_{t}=s_{t}$ when $s_{t}=1$; spatial size is $k_{s}=s_{s}$. Thus for the flexible layers the temporal kernel is 1 or 3 and the spatial kernel is 2 or 4. Spatial padding in these layers is “valid” (no padding).
Upsampling layers use causal $3\times 3\times 3$ convolutions. The convolution uses the same padding as elsewhere: causal in time (two leading frames, no trailing) and symmetric in space (one pixel on each side) so that spatial size is preserved before the depth-to-space rearrangement.

### 11 Mar 2026 v2

@@ 4.1 Key Findings
Our results provide the first systematic evidence on the role of tokeniser pretraining in physics foundation models. We show that pretraining substantially accelerates autoregressive training under limited computational budgets, with benefits critically dependent on domain alignment between pretraining data and autoregressive training data. Key findings include:
@@
Scale-dependent learning dynamics. Pretraining affects learning differently across spatial scales (Figure 2). At low frequencies, in-domain models achieve excellent final performance (NEPS $\sim$ 0.001) compared to models without pretraining (NEPS $\sim$ 0.02). At mid frequencies, in-domain models maintain consistently lower error throughout training (NEPS $\sim$0.1–0.3 vs. $\sim$0.3–1.0 without pretraining). At high frequencies, all models show poor quality (NEPS $\approx$ 1.0), but in-domain pretraining enables continuous improvement whereas other models show progressive degradation.
@@ 4.2 Limitations
Several limitations warrant discussion. First, out-of-domain pretraining provides modest benefits, and freezing the tokeniser in this setting degrades performance below the no-pretraining baseline. One explanation is limited field overlap between pretraining and downstream datasets: only one of the four physical quantities in the Euler dataset appears in the multi-dataset mixture (Table 1), while the others are encountered only during rollout training. Future work could investigate whether greater field overlap improves out-of-domain transfer, or develop methods to help pretrained tokenisers adapt to previously unseen physical fields.
@@ 4.3 Future Directions
Downstream task diversity. While this work focuses on autoregressive rollout prediction, pretrained tokenisers may prove valuable across a broader range of downstream settings. Applying the same pretrained tokeniser to diverse tasks and training objectives – such as masked token modelling, physical parameter inference, or inverse problems – would help reveal which settings benefit most from autoencoding pretraining.
Tokeniser architecture & training improvements. Investigating tokeniser design could yield more general insights and improve both pretraining and downstream accuracy. Promising directions include: evaluating alternative architectures to identify inductive biases better suited to physics data; scaling studies to determine whether benefits compound at larger capacity and with more diverse pretraining data; comparing training objectives (reconstruction, perceptual, adversarial losses) and quantisation strategies; and determining optimal pretraining budgets (measured in training steps, dataset size, or target accuracy) to maximise downstream performance while minimising computational cost.
Data composition & domain alignment. Our results show that domain alignment between pretraining and downstream datasets significantly impacts performance gains, raising an important question about data composition: given a fixed pretraining budget, does training on diverse multi-physics mixtures produce more versatile and generalisable tokenisers than domain-specific pretraining? Furthermore, systematically varying domain similarity, from small differences (boundary conditions, physical parameters) to large ones (governing equations), would quantify how domain proximity impacts transfer. Additionally, varying the size of the downstream dataset would reveal how pretraining gains scale with data availability and whether tokeniser pretraining is especially valuable in data-limited regimes.
@@
Together, these research directions would yield principled guidelines for tokeniser pretraining in physics foundation models, clarifying when pretraining justifies its computational cost and how data composition and domain alignment shape its benefits. Additionally, if future investigations demonstrate consistent out-of-domain benefits, they would justify developing and releasing high-quality pretrained tokenisers as community resources across scientific and engineering domains.
@@ 5 Conclusion
This work provides the first systematic investigation of tokeniser pretraining for physics foundation models. We demonstrate that pretraining offers a practical path to improved training efficiency, with benefits that depend critically on domain alignment between pretraining and downstream datasets. We also find that freezing most of an in-domain pretrained tokeniser can improve long-horizon rollouts, suggesting a simple regularisation strategy for stable autoregressive prediction. These findings position tokeniser pretraining as a simple yet effective mechanism for improving the training of physics foundation models.
@@ Appendix A Dataset Properties
Evolution of different gases starting with piecewise constant initial data in quadrants. The evolution can give rise to shocks, rarefaction waves, contact discontinuities, interaction with each other and domain walls. The evolved fields are density $\rho$, energy $e$, pressure $p$, and momentum components $(p_{x},p_{y})$, and the system follows Euler equations for a compressible gas. The data are sampled on a uniform $\times512$ Cartesian grid over 100 timesteps. The dataset comprises 20 parameter configurations, varying in boundary condition type (open or periodic) and gas constant $\gamma$, with 500 trajectories per configuration.
@@
Simulation of a continuum model for biological active matter, comprising rod-like particles immersed in a Stokes fluid. The evolved fields include a tracer-like concentration field $c$, fluid velocity $(v_{x},v_{y})$, an orientation tensor (describing alignment of the active agents), and a strain-rate tensor. The data are sampled on a uniform $\times256$ Cartesian grid over 81 timesteps, across 45 parameter sets and 5 trajectories per parameter set.
@@ Appendix B Tokeniser Architecture and Flexible Compression
MAGVIT-Simple autoencodes temporal sequences of 2D or 3D simulations ${\mathbf{x}}\in{\mathbb{R}}^{T\times S_{1}\times\dots\times S_{k}}$, where $k\in\{2,3\}$. The temporal size $T$ is set to “1 + a power of two”, and the spatial sizes $S_{i}$ are powers of two. Figure 4 illustrates the tokeniser architecture.
We employ causal convolutions, as introduced in MAGVIT-2, which enable joint image-video tokenisation and have been shown to outperform non-causal convolutions and attention-based blocks (Yu et al., 2024). For 2D simulations, a 3D causal convolutional layer with kernel size $(k_{t},k_{s},k_{s})$ is applied to the input. [^3] Operating on 3D simulations would require 4D causal convolutions. Spatial padding is applied symmetrically to both ends of each spatial dimension and is configured to preserve the input’s spatial size. Temporal padding is applied only at the beginning (i.e., $k_{t}-1$ leading frames), with no padding at the end. This preserves the temporal size of the input and ensures that each output timeframe depends only on the current and preceding frames. Padding values are set to zero.
The encoder and decoder heads use adaptive causal 3D convolutions that support a variable number of input physical fields. A single convolutional kernel is learned for the maximum number of fields, $c_{\text{total}}$. At runtime, given an input with $c_{\text{active}}$ fields, rows and columns corresponding to active fields are selected from the weight matrix, and the output is rescaled by $\sqrt{c_{\text{total}}/c_{\text{active}}}$ to compensate for reduced fan-in. This follows the field embedding approach from MPP (McCabe et al., 2024).
@@
Given a scaling pair $(s_{t},s_{s})$, we define the spatial and temporal kernel sizes as:
$k_{s}=s_{s}\qquad\qquad\qquad k_{t}=\begin{cases}1+s_{t}&\text{if }s_{t}>1,\\ s_{t}&\text{if }s_{t}=1.\end{cases}\;.$ (6)
@@
pixel in the output; a strategy we adopt in our flexible depth-to-space layers. Finally, because this operation introduces additional timeframes, we follow MAGVIT-2 (Yu et al., 2024) and discard the first $s_{t}-1$ frames from the output.
@@ Appendix C Detailed Training Recipe
This appendix provides the complete set of hyperparameters for both training stages. High-level training details are described in Subsection 2.5, and the tokeniser architecture is described in Appendix B. Table 6 summarises the training hyperparameters.
@@
For rollout training, we use an inverse-square-root schedule with linear warmup and square-root cooldown, stepping once per epoch. Let $e$ denote the current epoch (0-indexed), with total epochs $E$, warmup duration $W$, cooldown duration $C$, warmup starting factor $\alpha=0.1$, and decay offset $\kappa=128$. The schedule has three phases:
@@
$\text{lr}(e)=\text{lr}_{\text{peak}}\left(\alpha+(1-\alpha)\frac{e}{W}\right).$ (7)
@@
$\text{lr}(e)=\frac{\text{lr}_{\text{peak}}}{1+\sqrt{s+\kappa}-\sqrt{\kappa}}\,$, (8)
where s:=eWs:=e-W counts epochs into decay ($s=1,2,\ldots$).
Cooldown (e>ECe>E-C): The learning rate ramps down with a square-root profile:
$\text{lr}(e)=\text{lr}_{\text{end}}\left(1-\sqrt{\frac{t-1}{C}}\right),$ (9)
where t:=e(EC)t:=e-(E-C) counts epochs into cooldown ($t=1,2,\ldots$) and
$\text{lr}_{\text{end}}:=\frac{\text{lr}_{\text{peak}}}{1+\sqrt{E-C+\kappa}-\sqrt{\kappa}}\,.$ (10)
@@
Each GPU processes a disjoint shard of the training data. For tokeniser pretraining, each shard contains 21,000 unique batches (of size 2). Once all unique batches have been consumed, the data loader reseeds its random number generator and begins iterating over the dataset again, allowing resampling in subsequent passes. For rollout training, the same mechanism applies but each shard contains 2100 unique batches.
@@
In downsampling layers, stride equals the chosen compression $(s_{t},s_{s},s_{s})$ (see Table 7). The convolution kernel size is tied to the stride: temporal size is $k_{t}=1+s_{t}$ when $s_{t}>1$ and $k_{t}=s_{t}$ when $s_{t}=1$; spatial size is $k_{s}=s_{s}$. Thus for the flexible layers the temporal kernel size is 1 or 3 and the spatial kernel size is 2 or 4. Spatial padding in these layers is “valid” (no padding).
We use depth-to-space upsampling. The corresponding layers employ causal $3\times 3\times 3$ convolutions. These convolutions follow the same padding scheme as elsewhere: causal in time (two leading frames, no trailing) and symmetric in space (one pixel on each side) so that spatial size is preserved before the depth-to-space rearrangement.